OpenStudy (anonymous):

What is the solution to the inequality -50n is greater than or equal to 10

OpenStudy (goldphenoix):

\[\large \large -50m \ge 10\] You want to find m, so we want to get rid of -50. -50m is a coefficient. So another way of writing that is -50 * m. To get rid of -50, we want to do the reciprocal of multiplication. What is the reciprocal of multiplication?

OpenStudy (anonymous):

n is less then or equal to -1/5

OpenStudy (anonymous):

Two different answers? And to reword the question its 

OpenStudy (texaschic101):

-50n <= 10 This problem says less then or equal....is this the correct problem

OpenStudy (anonymous):

Yes \[-50n \le 10\]

OpenStudy (goldphenoix):

What do you mean, Baby?

OpenStudy (anonymous):

−50n≤10 is the question i need help with

OpenStudy (texaschic101):

inequalities are solved like equations. You want to get n on one side and everything else on the other side. One rule for inequalities that you have to remember is when dividing by a negative, he inequality sign changes. I will let GoldPhoenix finish because he started working with you.

OpenStudy (goldphenoix):

~Total confusion~ So the question you posted is different from the question you need help with?

OpenStudy (anonymous):

I just noticed i could use the equation button and i ment less than or equal two in my main question sorry :/

OpenStudy (goldphenoix):

Ah. Okay. So you would have: \[\large \large -50n \le 10\] Now you remember what I said before, correct? So what is the reciprocal of multiplication?

OpenStudy (goldphenoix):

Or if you don't understand what I mean, then how can we get n by itself?

OpenStudy (anonymous):

Division :D

OpenStudy (goldphenoix):

Great! So what number divided by -50 will give us 1?

OpenStudy (anonymous):

-50

OpenStudy (goldphenoix):

Good job. So -50 / -50 = 1 If you do that to the left side, then you must do it to the right side. What does 10 / -50 give us?

OpenStudy (anonymous):

Like -0.2 or somthing like that

OpenStudy (goldphenoix):

Great. Now here's the thing. We must switch the inequality because you divided both side by a negative number. So what's the opposite of less than or equal to?

OpenStudy (goldphenoix):

The only time when we are switching the inequality sign around is when we are either multiplying or dividing both side by a negative number.

OpenStudy (anonymous):

Lol greater than or equal to

OpenStudy (goldphenoix):

Great. So your final answer should be: \[\large \large n \ge -0.2\]

OpenStudy (anonymous):

Herm. Oh yes! It just has it in division form. Thank you very much